OpenStudy (anonymous):

Solve by substitution. x=5y+21 x=5/4y (it's a fraction)

OpenStudy (raffle_snaffle):

set them equal to each other and solve for y

OpenStudy (anonymous):

set them equal how? sorry i'm really bad with these.

OpenStudy (shamil98):

is this (5/4)y or (5/4y) ?

OpenStudy (anonymous):

(5/4)

OpenStudy (anonymous):

oh wait 5/4y sorry

OpenStudy (anonymous):



OpenStudy (shamil98):

kk. \[\large \frak \frac{ 5 }{ 4 }y = 5y + 21\] Does this help?

OpenStudy (anonymous):

a little, the fractions just make it all the more confusing haha.

OpenStudy (shamil98):

Just solve for y now \[\huge \frak \frac{ -15 }{ 4 }y= 21\] Multiply by the reciprocal. \[\huge \frak y = \frac{ -84 }{ 15 } = \frac{ -28 }{ 5 }\]

OpenStudy (shamil98):

\[\huge \frak x = \frac{ 5 }{ 4 }y\] \[\huge \frak x = \frac{ 5 }{ 4 } (\frac{-28}{5})\] \[\huge \frak x = -7\]

OpenStudy (anonymous):

okay so then you incorporate -7 into the equation? right?

OpenStudy (shamil98):

No, you already have your x and y values.

OpenStudy (shamil98):

x = -7 y = -28/5

OpenStudy (anonymous):

ohhh okay i see.

OpenStudy (anonymous):

thanks :)